What velocity must a 15kg object be moving at in order for it to have a momentum of 54kgm/s?
1
Expert's answer
2020-12-10T11:02:18-0500

By the definition of the momentum, we have:


"p=mv,""v=\dfrac{p}{m}=\dfrac{54\ \dfrac{kgm}{s}}{15\ kg}=3.6\ \dfrac{m}{s}."

Answer:

"v=3.6\ \dfrac{m}{s}."